OpenStudy (anonymous):

Find the slope and y-intercept of the line 4x + 5y − 1 = 0

hero (hero):

If you write it in the form y = mx + b, m = slope and b = y-intercept

OpenStudy (anonymous):

basically solve the equation for y: \[y = -\frac{4}{5}x + \frac{1}{5}\] now you can tell what m and b are....what are they?

OpenStudy (anonymous):

m=-4/5 b=1/5

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

whats my y intercept?

OpenStudy (anonymous):

y intercept is the number which is floating around once you have the equation in the form y=mx + b. It is b. Which means in our example the y-intercept is 1/5

OpenStudy (anonymous):

and my x would be?

OpenStudy (anonymous):

x is allowed to vary across the interval on which the line is defined, which is usually the entire number line.

OpenStudy (anonymous):

your gradient is m, and in this case it is -4/5.

OpenStudy (anonymous):

do you see that as x varies y varies too?

OpenStudy (anonymous):

ok. im confused how i would graph the x,y intercept

OpenStudy (anonymous):

put a pencil dot on the y-axis at 1/5. Count -4 units down on the y axis, then count 5 units to the right and then put a pencil dot there. Connect the two dots and that is your line.

OpenStudy (anonymous):

yes that picture is good. So you start at 1/5 on the y-axis and make a mark. Then do 1/5 - 4 = -19/5. Then at -19/5 on the y-axis you count out 5 units to the right and make a mark. Those two marks are joined to make your line.

OpenStudy (anonymous):

ugh! my x would be 5 or 4?

OpenStudy (anonymous):

x has no value unless you give y a value...and conversely.

OpenStudy (anonymous):

ok. still not getting it. thNKS